OpenStudy (anonymous):

Solve the following for x. x divided by ten is greater than or equal to fifteen x is greater than or equal to fifteen. x equals one hundred fifty x is greater than one hundred fifty x is greater than or equal to one hundred fifty

OpenStudy (anonymous):

\[\frac{ x }{ 10 } \ge 15\] simplifies to: \[x \ge 150\] all of the rest are already reduced to x.